# Universal relations for fundamental modes of rotating neutron stars with differential rotations

### v1

@@ 4 Fundamental mode frequencies against kinetic-to-binding energy ratio
As mentioned in Section 1, through simulations with the Cowling approximation, Stergioulas et al. (2004) have demonstrated that the frequency of axisymmetric oscillation modes scales almost linearly with the kinetic-to-binding energy ratio for rotating neutron stars. By assuming a fixed relative differences between the actual frequencies and those obtained in their simulations with the Cowling approximation, they have also constructed linear relations as functions of $T/|W|$ for predicting fundamental $l=0$ quasi-radial mode frequency $f_{F}$ and fundamental $l=2$ quadrupolar mode frequency $f_{{}^{2}f}$. Nonetheless, Dimmelmeier et al. (2006) later showed that these linear relations can contribute to an overestimate of $f_{F}$ and underestimate of $f_{{}^{2}f}$ up to $\sim 30\%$. In addition, both studies only regarded differentially rotating neutron stars with a single value of degree of differential rotation $\tilde{A}=1$. Hence, we revisit the relation between the fundamental mode frequency $f$ and kinetic-to-binding energy ratio $T/|W|$ by considering both the effect of dynamical spacetime and various degrees of differential rotation. We plot fundamental $l=0$ quasi-radial mode frequency $f_{F}$ (top left panel) and fundamental $l=2$ quadrupolar mode frequency $f_{{}^{2}f}$ (top right panel) against kinetic-to-binding energy ratio $T/|W|$ in Fig. 2. The data points are arranged into 5 sequences with $\tilde{A}\in\{0.0,1.0,2.0,3.0,4.0\}$, where $\tilde{A}$ is the degree of differential rotation. $\tilde{A}=0.0$ refers to uniformly rotating cases. We observe that the frequencies $f$ of both fundamental modes for models with different $\tilde{A}$ decrease roughly linearly with $T/|W|$.
@@
We compute the deviations between our simulation data $f^{\mathrm{data}}$ and our linear fits $f^{\mathrm{pred}}$. We then plot the deviations for $F$-mode $f^{\mathrm{data}}_{F}/f^{\mathrm{pred}}_{F}-1$ (bottom left panel) and ${}^{2}f$-mode $f^{\mathrm{data}}_{{}^{2}f}/f^{\mathrm{pred}}_{{}^{2}f}-1$ (bottom right panel) against $T/|W|$ in Fig. 2. We find that only a slight deviation between our simulation data and our linear fits with $f^{\mathrm{data}}_{F}/f^{\mathrm{pred}}_{F}-1\lesssim 1\%$ and $f^{\mathrm{data}}_{{}^{2}f}/f^{\mathrm{pred}}_{{}^{2}f}-1\lesssim 2\%$. We also compare our linear fits with the data points of the previous study of Dimmelmeier et al. (2006) (purple diamonds), in which the fundamental modes of uniformly rotating and differentially rotating ($\tilde{A}=1.0$) neutron stars in dynamical spacetime were considered. We find that only minor deviations between the simulation data in Dimmelmeier et al. (2006) and our linear fits with $f^{\mathrm{data}}/f^{\mathrm{pred}}-1\lesssim 2\%$ for both fundamental modes. Thus, we have constructed universal relations that can quantify $f_{F}$ and $f_{{}^{2}f}$ of rotating neutron stars with differential rotations in dynamical spacetime.
We compare our universal relations with those constructed in the previous work with the Cowling approximation by Stergioulas et al. (2004). For the predicted quasi-radial $l=0$ fundamental mode frequency $f_{F}$, they proposed
@@
These deviations are mainly due to the assumption of a fixed relative differences between the actual frequencies and those obtained in their simulation with the Cowling approximation. In Stergioulas et al. (2004), they adopted an empirical relation between the actual frequency and the frequency in the Cowling approximation proposed by Font et al. (2002) for constructing their linear fits,
@@
where the superscript (C) denotes the quantity in the Cowling approximation, $f_{\Omega}$ is the frequency of the fundamental mode for a rotating star with angular velocity $\Omega$ , and $f_{0}$ is the corresponding frequency in a non-rotating star. This empirical relation was initially proposed based on uniformly rotating cases. Hence, they assumed a fixed relative difference $f_{0}-f_{0}^{(\mathrm{C})}$ for all models. As discussed in Dimmelmeier et al. (2006), the relative differences between the actual frequencies and those obtained in the Cowling approximation are significantly changed as the kinetic-to-binding energy ratio $T/|W|$ of the star increases. Therefore, this assumption of a fixed relative difference would cause a deviation between our linear fits and those obtained in Dimmelmeier et al. (2006).
@@
Figure 2: Plots of fundamental mode frequencies $f$ (top panels) and the frequency deviations between the simulation data $f^{\mathrm{data}}$ and the predictions by linear fits $f^{\mathrm{pred}}$ (bottom panels) against kinetic-to-binding energy ratio $T/|W|$. In particular, we plot fundamental $l=0$ quasi-radial mode frequency $f_{F}$ (top left panel), deviation of $f_{F}$ (bottom left panel), fundamental $l=2$ quadrupolar mode frequency $f_{{}^{2}f}$ (top right panel), and deviation of $f_{{}^{2}f}$ (bottom right panel) against $T/|W|$. The data points are arranged into 5 sequences with $\tilde{A}\in\{0.0,1.0,2.0,3.0,4.0\}$, where $\tilde{A}$ is the degree of differential rotation. $\tilde{A}=0.0$ refers to uniformly rotating models. Both $f_{F}$ and $f_{{}^{2}f}$ for different $\tilde{A}$ decrease almost linearly with $T/|W|$. Hence, we perform linear regressions using our simulation data to obtain 2 linear relations of $f^{\mathrm{pred}}_{F}(T/|W|)$ in Eq. (6) and $f^{\mathrm{pred}}_{{}^{2}f}(T/|W|)$ in Eq. (7) respectively. We find that only a slight deviation between our simulation data and our linear fits with $f^{\mathrm{data}}_{F}/f^{\mathrm{pred}}_{F}-1\lesssim 1\%$ and $f^{\mathrm{data}}_{{}^{2}f}/f^{\mathrm{pred}}_{{}^{2}f}-1\lesssim 2\%$. We also compare our linear fits with the data points of the previous study of Dimmelmeier et al. (2006) (purple diamonds), in which the fundamental modes of uniformly rotating and differentially rotating ($\tilde{A}=1.0$) neutron stars in dynamical spacetime were considered. We find that only minor deviations between the simulation data in Dimmelmeier et al. (2006) and our linear fits with $f^{\mathrm{data}}/f^{\mathrm{pred}}-1\lesssim 2\%$ for both fundamental modes. Therefore, we have constructed universal relations that can quantify $f_{F}$ and $f_{{}^{2}f}$ for rotating neutron stars with differential rotations in dynamical spacetime.
5 Conclusions
@@
In this work, for the first time, we considered various degrees of differential rotation $\tilde{A}$ and dynamical spacetime to construct linear relations quantifying the frequencies of fundamental $l=0$ quasi-radial mode $f_{F}$ and fundamental $l=2$ quadrupolar mode $f_{{}^{2}f}$. In particular, we first found that fundamental mode frequency $f$ scales nearly linearly with stellar compactness $M/R$. Next, we observed quasi-linear relations between $f$ and the ratio of kinetic energy to binding energy $T/|W|$ for different values of $\tilde{A}$. Lastly, we constructed linear relations that can quantify $f_{F}$ and $f_{{}^{2}f}$ of rotating neutron stars with differential rotations based on linear regressions. In consequence, these relations can serve as universal relations that can quantify $f_{F}$ and $f_{{}^{2}f}$ for rotating neutron stars with differential rotations in dynamical spacetime..
@@
Several natural extensions should be considered to model differentially rotating neutron stars in different scenarios more realistically. First, a more realistic equation of state containing thermal effects should be used. In addition, ultra-strong magnetic field strength $B\sim\mathcal{O}(10^{17})$ G could be generated during the formation of a proto-neutron stars (Pili et al., 2014), and binary neutron star mergers (Price & Rosswog, 2006). This ultra-strong magnetic field strength could also affect the frequency of fundamental modes (Leung et al., 2022). Thus, the effect of the magnetic field should also be considered. Also, more realistic rotational profiles (e.g. Uryū et al. 2017) could be adopted to model differentially rotating neutron stars in different astrophysical scenarios. Lastly, 3D simulations without axisymmetry should be conducted to examine the non-axisymmetric oscillation modes.
@@ Appendix A Equilibrium models of neutron stars
This appendix summarizes the detailed properties of the equilibrium models of the 5 sequences in this study. These data were used to construct the linear fits discussed in the main text. We provide the data in the form of one table for each sequence. Each sequence refers to each value of the degree of differential rotation $\tilde{A}\in\{0.0,1.0,2.0,3.0,4.0\}$, where $\tilde{A}=0.0$ refers to uniform rotation. In each table, we list the model, the central rest mass density $\rho_{\mathrm{c}}$, the gravitational mass $M$, the circumferential radius $R$, the equatorial radius $r_{\mathrm{e}}$, the ratio of polar radius to equatorial radius $r_{\mathrm{p}}/r_{\mathrm{e}}$, the central angular velocity $\Omega_{\mathrm{c}}$, the ratio of central angular velocity to equatorial angular velocity $\Omega_{\mathrm{c}}/\Omega_{\mathrm{e}}$, and the kinetic-to-binding energy ratio $T/|W|$.

### v2

@@ IV Fundamental mode frequencies against kinetic-to-binding energy ratio
As mentioned in Section I, through simulations with the Cowling approximation, [6] have demonstrated that the frequency of axisymmetric oscillation modes scales almost linearly with the kinetic-to-binding energy ratio for rotating neutron stars. By assuming a fixed relative differences between the actual frequencies and those obtained in their simulations with the Cowling approximation, they have also constructed linear relations as functions of $T/|W|$ for predicting fundamental $l=0$ quasi-radial mode frequency $f_{F}$ and fundamental $l=2$ quadrupolar mode frequency $f_{{}^{2}f}$. Nonetheless, [17] later showed that these linear relations can contribute to an overestimate of $f_{F}$ and underestimate of $f_{{}^{2}f}$ up to $\sim 30\%$. In addition, both studies only regarded differentially rotating neutron stars with a single value of degree of differential rotation $\tilde{A}=1$. Hence, we revisit the relation between the fundamental mode frequency $f$ and kinetic-to-binding energy ratio $T/|W|$ by considering both the effect of dynamical spacetime and various degrees of differential rotation. We plot fundamental $l=0$ quasi-radial mode frequency $f_{F}$ (top left panel) and fundamental $l=2$ quadrupolar mode frequency $f_{{}^{2}f}$ (top right panel) against kinetic-to-binding energy ratio $T/|W|$ in Fig. 2. The data points are arranged into 5 sequences with $\tilde{A}\in\{0.0,1.0,2.0,3.0,4.0\}$, where $\tilde{A}$ is the degree of differential rotation. $\tilde{A}=0.0$ refers to uniformly rotating cases. We observe that the frequencies $f$ of both fundamental modes for models with different $\tilde{A}$ decrease roughly linearly with $T/|W|$.
@@
We compute the deviations between our simulation data $f^{\mathrm{data}}$ and our linear fits $f^{\mathrm{pred}}$. We then plot the deviations for $F$-mode $f^{\mathrm{data}}_{F}/f^{\mathrm{pred}}_{F}-1$ (bottom left panel) and $^{2}f-$mode $f^{\mathrm{data}}_{{}^{2}f}/f^{\mathrm{pred}}_{{}^{2}f}-1$ (bottom right panel) against $T/|W|$ in Fig. 2. We find that only a slight deviation between our simulation data and our linear fits with $f^{\mathrm{data}}_{F}/f^{\mathrm{pred}}_{F}-1\lesssim 1\%$ and $f^{\mathrm{data}}_{{}^{2}f}/f^{\mathrm{pred}}_{{}^{2}f}-1\lesssim 2\%$. We also compare our linear fits with the data points of the previous study of [17] (purple diamonds), in which the fundamental modes of uniformly rotating and differentially rotating ($\tilde{A}=1.0$) neutron stars in dynamical spacetime were considered. We find that only minor deviations between the simulation data in [17] and our linear fits with $f^{\mathrm{data}}/f^{\mathrm{pred}}-1\lesssim 2\%$ for both fundamental modes. Thus, we have constructed linear relations that can quantify $f_{F}$ and $f_{{}^{2}f}$ of rotating neutron stars with differential rotations in dynamical spacetime.
We compare our linear relations with those constructed in the previous work with the Cowling approximation by [6]. For the predicted quasi-radial $l=0$ fundamental mode frequency $f_{F}$, they proposed
@@
These deviations are mainly due to the assumption of a fixed relative differences between the actual frequencies and those obtained in their simulation with the Cowling approximation. In [6], they adopted an empirical relation between the actual frequency and the frequency in the Cowling approximation proposed by [14] for constructing their linear fits,
@@
where the superscript (C) denotes the quantity in the Cowling approximation, $f_{\Omega}$ is the frequency of the fundamental mode for a rotating star with angular velocity $\Omega$ , and $f_{0}$ is the corresponding frequency in a non-rotating star. This empirical relation was initially proposed based on uniformly rotating cases. Hence, they assumed a fixed relative difference $f_{0}-f_{0}^{(\mathrm{C})}$ for all models. As discussed in [17], the relative differences between the actual frequencies and those obtained in the Cowling approximation are significantly changed as the kinetic-to-binding energy ratio $T/|W|$ of the star increases. Therefore, this assumption of a fixed relative difference would cause a deviation between our linear fits and those obtained in [17].
@@
Figure 2: Plots of fundamental mode frequencies $f$ (top panels) and the frequency deviations between the simulation data $f^{\mathrm{data}}$ and the predictions by linear fits $f^{\mathrm{pred}}$ (bottom panels) against kinetic-to-binding energy ratio $T/|W|$. In particular, we plot fundamental $l=0$ quasi-radial mode frequency $f_{F}$ (top left panel), deviation of $f_{F}$ (bottom left panel), fundamental $l=2$ quadrupolar mode frequency $f_{{}^{2}f}$ (top right panel), and deviation of $f_{{}^{2}f}$ (bottom right panel) against $T/|W|$. The data points are arranged into 5 sequences with $\tilde{A}\in\{0.0,1.0,2.0,3.0,4.0\}$, where $\tilde{A}$ is the degree of differential rotation. $\tilde{A}=0.0$ refers to uniformly rotating models. Both $f_{F}$ and $f_{{}^{2}f}$ for different $\tilde{A}$ decrease almost linearly with $T/|W|$. Hence, we perform linear regressions using our simulation data to obtain 2 linear relations of $f^{\mathrm{pred}}_{F}(T/|W|)$ in Eq. (6) and $f^{\mathrm{pred}}_{{}^{2}f}(T/|W|)$ in Eq. (7) respectively. We find that only a slight deviation between our simulation data and our linear fits with $f^{\mathrm{data}}_{F}/f^{\mathrm{pred}}_{F}-1\lesssim 1\%$ and $f^{\mathrm{data}}_{{}^{2}f}/f^{\mathrm{pred}}_{{}^{2}f}-1\lesssim 2\%$. We also compare our linear fits with the data points of the previous study of [17] (purple diamonds), in which the fundamental modes of uniformly rotating and differentially rotating ($\tilde{A}=1.0$) neutron stars in dynamical spacetime were considered. We find that only minor deviations between the simulation data in [17] and our linear fits with $f^{\mathrm{data}}/f^{\mathrm{pred}}-1\lesssim 2\%$ for both fundamental modes. Therefore, we have constructed linear relations that can quantify $f_{F}$ and $f_{{}^{2}f}$ for rotating neutron stars with differential rotations in dynamical spacetime.
V Conclusions
@@
In this work, for the first time, we considered various degrees of differential rotation $\tilde{A}$ and dynamical spacetime to construct linear relations quantifying the frequencies of fundamental $l=0$ quasi-radial mode $f_{F}$ and fundamental $l=2$ quadrupolar mode $f_{{}^{2}f}$. In particular, we first found that fundamental mode frequency $f$ scales nearly linearly with stellar compactness $M/R$. Next, we observed quasi-linear relations between $f$ and the ratio of kinetic energy to binding energy $T/|W|$ for different values of $\tilde{A}$. Lastly, we constructed linear relations that can quantify $f_{F}$ and $f_{{}^{2}f}$ of rotating neutron stars with differential rotations based on linear regressions. In consequence, these relations can quantify $f_{F}$ and $f_{{}^{2}f}$ for rotating neutron stars with differential rotations in dynamical spacetime.
@@
Several natural extensions should be considered to model differentially rotating neutron stars in different scenarios more realistically. First, for a more accurate determination of mode frequencies in neutron stars, particularly in scenarios such as core-collapse supernovae and neutron star mergers, it is important to use a more realistic equation of state that includes thermal effects (see e.g. [3, 40]) as well as empirically motivated rotational laws (see e.g. [36, 37, 38, 39, 41, 42]) as proposed in the literature for modeling these scenarios. In addition, 3D simulations that do not assume axisymmetry should be performed to examine non-axisymmetric oscillation modes. These modes can arise and become significant in binary neutron star remnants and proto-neutron stars formed during core-collapse supernovae (see e.g. [25]). Furthermore, ultra-strong magnetic fields with strengths $B\sim\mathcal{O}(10^{17})\mathrm{G}$ can be generated during the formation of proto-neutron stars [32] and in binary neutron star mergers [55]. Such strong magnetic fields may also affect the frequencies of fundamental modes [27]. Therefore, the effects of magnetic fields should also be taken into account.
Acknowledgements.
@@ Appendix A Equilibrium models of neutron stars
This appendix summarizes the detailed properties of the equilibrium models of the 5 sequences in this study. These data were used to construct the linear fits discussed in the main text. We provide the data in the form of one table for each sequence (see Table A1 - A5). Each sequence refers to each value of the degree of differential rotation $\tilde{A}\in\{0.0,1.0,2.0,3.0,4.0\}$, where $\tilde{A}=0.0$ refers to uniform rotation. In each table, we list the model, the central rest mass density $\rho_{\mathrm{c}}$, the gravitational mass $M$, the circumferential radius $R$, the equatorial radius $r_{\mathrm{e}}$, the ratio of polar radius to equatorial radius $r_{\mathrm{p}}/r_{\mathrm{e}}$, the central angular velocity $\Omega_{\mathrm{c}}$, the ratio of central angular velocity to equatorial angular velocity $\Omega_{\mathrm{c}}/\Omega_{\mathrm{e}}$, and the kinetic-to-binding energy ratio $T/|W|$.
@@
Appendix B Mode extraction and identification
To illustrate how we identify and extract the fundamental modes, we plot the power spectral density (PSD) of the $r$-component $v_{r}$ with an $l=0$ perturbation imposed (left panels) and the $\theta$-component $v_{\theta}$ with an $l=2$ perturbation imposed (right panels) of the three-velocity field in arbitrary units for the non-rotating model A0 (top panels) and model E25 (bottom panels) in Fig.A1. Since our non-rotating model A0 corresponds to ‘BU0’ (or ‘AU0’) in the previous study by Dimmelmeier et al.[17], we can identify the peaks in the PSDs of $v_{r}$ and $v_{\theta}$ that correspond to the fundamental $l=0$ quasi-radial mode frequency $f_{F}$ and the fundamental $l=2$ quadrupolar mode frequency $f_{{}^{2}f}$ for our model A0 by comparing them with the well-tested mode frequencies reported by Dimmelmeier et al. (dotted lines), as shown in the top panels of Fig.A1. After identifying the peaks of $f_{F}$ and $f_{{}^{2}f}$ in the PSDs, we track how $f_{F}$ and $f_{{}^{2}f}$ change in each sequence as the kinetic-to-binding energy ratio $T/|W|$ increases. As shown in the bottom panels of Fig.A1, we take model E25, which is one of the models in the sequence with the highest degree of differential rotation ($\tilde{A}=4.0$), as an illustrative example to demonstrate how we extract $f_{F}$ and $f_{{}^{2}f}$ (dashed lines) from the PSDs.
Figure A1: The power spectral density (PSD) of the $r$-component $v_{r}$ with $l=0$ perturbation imposed (left panels) and the $\theta$-component $v_{\theta}$ with $l=2$ perturbation imposed (right panels) of the three-velocity field in arbitrary units for the non-rotating model A0 (top panels) and model E25 (bottom panels). We first identify the peaks in the PSDs that correspond to the fundamental $l=0$ quasi-radial mode frequency $f_{F}$ and the fundamental $l=2$ quadrupolar mode frequency $f_{{}^{2}f}$ for our model A0 (top panels) by comparing them with the well-tested mode frequencies reported by Dimmelmeier et al. (dotted lines). After identifying the peaks of $f_{F}$ and $f_{{}^{2}f}$ in the PSDs, we track how $f_{F}$ and $f_{{}^{2}f}$ change in each sequence as the kinetic-to-binding energy ratio $T/|W|$ increases, as illustrated in the example of model E25 (bottom panels).